OpenStudy (anonymous):

I need to solve for x in 14-(6x-5)=5(2x-1)-4x!! Help please and explain thoroughly, thank you

OpenStudy (anonymous):

14-(6x-5)=5(2x-1)-4x 14-6x+5=10x-5-4x 19-6x=6x-5 24=12x x=2

OpenStudy (anonymous):

i think that explains enough :)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

It's a little confusing though, I did this. I'm not sure if it's the same as ur demonstration:

OpenStudy (anonymous):